OpenStudy (anonymous):

Write each equation in standard form using integers. y+4=1/3(x+6)

OpenStudy (anonymous):

y+4=1/3(x+6) y+4=1/3(x) + 1/3(6) y+4=1/3x + 2 now, substract 4 from boths sides y+4-4=1/3x+2-4 y=1/3x-2

OpenStudy (anonymous):

standard form: ax+by=c 1/3x+y=-2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i forgot to include the - sign. -1/3x+y=-2 final answer.